my daughter is struggling with this maths she is 9 yr,, i do not know how to give the answer or explain how to do it.. please can you help me (mum)..

The Highwayman steals 25000kg of gold coins. Later he steals 3kg of gold coins. He needs 31kg of gold to buy himself a coach to rife around in how many more grams of gold does he need?
1
Expert's answer
2011-11-22T09:35:13-0500
if the coins are gold, i.e.& that gay stole& 25 000 kg of golg. latert he got 3kg more. hence he has 25000+3=25003 kg of gold. if he need only 31 kg he already has it
if you& made a mistake and he stole 25 kg(not 25000kg) of coins and later 3 then he has 25+3=28 kg. If he needs 31 kg he has to get 31-28=3 kg=3000 g& morehope it will help